OpenStudy (anonymous):

please solve 4/7x=8/35

OpenStudy (katrinakaif):

Mulitply by the reciprocal 4/7x = 8/35 x = 8/35 x 7/4 And you get...?

OpenStudy (anonymous):

8/20

OpenStudy (karatechopper):

0.4 or 8/20

OpenStudy (katrinakaif):

Now simplify 8/20 It is divisible by 4

OpenStudy (anonymous):

Thank you!

OpenStudy (katrinakaif):

8/20 = 2/5 would be your final answer Anytime~ =)

OpenStudy (anonymous):

:)